### v1

# Higher order linear differential equations for unitary matrix integrals: applications and generalisations

Abstract
We consider characterisations of the class of unitary matrix integrals $\langle(\det U)^{q}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\rangle_{U(l)}$ in terms of a first order matrix linear differential equation for a vector function of size $l+1$, and in terms of a scalar linear differential equation of degree $l+1$. It will be shown that the latter follows from the former. The matrix linear differential equation provides an efficient way to compute the power series expansion of the matrix integrals, which with $q=0$ and $q=l$ are of relevance to the enumeration of longest increasing subsequences for random permutations, and to the question of the moments of the first and second derivative of the Riemann zeta on the critical line respectively. This procedure is compared against that following from known characterisations involving the $\sigma$-Painlevé III second order nonlinear differential equation. We show too that that the natural $\beta$ generalisation of the unitary group integral permits characterisation by the same classes of linear differential equations.
1School of Mathematics and Statistics, The University of Melbourne, Victoria 3010, Australia.     Email: pjforr@unimelb.edu.au
2Department of Mathematics, University of Sussex, Brighton, BN1 9RH, United Kingdom.
Email: weif0831@gmail.com
1 Introduction
1.1 A unitary matrix integral and the longest increasing subsequence problem
The study of this article is the systematic derivation of linear differential equations satisfied by
$\Big{\langle}e^{z{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)},$ (1.1)
the same average with a further factor of $(\det U)^{q}$, and their $\beta$ generalisation. In (1.1) $U(l)$ refers to the group of $l\times l$ unitary matrices sampled with Haar measure. This matrix integral first appeared in the theoretical physics literature on two-dimensional Yang-Mills gauge theory as a partition function [39]. However it is not in this context that interest in differential equations for matrix integrals first arose. Rather, it came about due to the realisation that the coefficients $\{T_{l}(N)\}$ say of $z^{2N}/(N!)^{2}$ in the power series of (1.1) (note that this is even in $z$) are for a given $l$ an integer family, e.g. for $l=4$ and starting at $N=1$ [40, sequence A047889]
$1,2,6,24,119,694,4582,33324,261808,2190688,19318688,178108704,\dots$ (1.2)
Thus, according to [42], $T_{l}(N)$ counts the number or permutations of $\{1,2,\dots,N\}$ such that the length of the longest increasing subsequence is less than or equal to $l$ (this number is $N!$ for $N\leq l$ as illustrated in (1.2)).
In a different guise a formula equivalent to
$\Big{\langle}e^{z{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}z^{2N}$ (1.3)
appeared in the earlier work [32]. There it was shown
$\det\Big{[}I_{j-k}(2z)\Big{]}_{j,k=1,\dots,l}=1+\sum_{N=1}^{\infty}{T_{l}(N)\over(N!)^{2}}z^{2N},$ (1.4)
where $I_{\nu}(x)={1\over 2\pi}\int_{0}^{2\pi}e^{x\cos\theta}e^{i\nu\theta}\,d\theta$, $\nu\in\mathbb{Z}$, is the modified Bessel function of the first kind (i.e. I-Bessel function), although $T_{l}(N)$ was defined as the number of pairs of standard tableaus of content $N$ with length of the first row less than or equal to $l$. For an explanation of these notions, and the equivalence to increasing subsequences of permutations, see e.g. [31, 43]. The reason why the Toeplitz determinant in (1.4) is equivalent to the matrix integral in (1.3) follows from the general identities
$\Big{\langle}\prod_{j=1}^{l}a(\theta_{j})\Big{\rangle}_{U(l)}={1\over(2\pi)^{l}l!}\int_{0}^{2\pi}d\theta_{1}\cdots\int_{0}^{2\pi}d\theta_{l}\,\prod_{j=1}^{l}a(\theta_{j})\prod_{1\leq j<k\leq l}|e^{i\theta_{k}}-e^{i\theta_{j}}|^{2}\\ =\det\Big{[}{1\over 2\pi}\int_{0}^{2\pi}a(\theta)e^{i(j-k)\theta}\,d\theta\Big{]}_{j,k=1,\dots,l}.$ (1.5)
Here the first equality follows from Weyl’s result on the explicit functional form of the eigenvalue probability density function (PDF) for matrices from $U(l)$ with Haar measure [46], [18], while the second may be regarded as a circular analogue of Andréief’s identity for the integral over the product of two determinants [22], with use too of the Vandermonde determinant formula to rewrite the product over the differences squared in this form.
Nonconstructive arguments were given in [32] that the generating function in (1.3) is $D$-finite, meaning that it must satisfy a linear differential equation with polynomial coefficients. Later, in [9], these differential equations where determined explicitly for $l\leq 7$. For example, in the case $l=4$ and thus relating to (1.2) one has that (1.3) with $z\mapsto\sqrt{z}$ is a solution of the fifth order equation [9, last collection of displayed equations]
$z^{4}Y^{(5)}+20z^{3}Y^{(4)}-2(10z-59)z^{2}Y^{(3)}-2(91z-110)zY^{\prime\prime}+4(16z^{2}-87z+20)Y^{\prime}+16(8z-5)Y=0$ (1.6)
(there is also a known linear differential equation of degree four for the power series formed according to the rule that the coefficient of $z^{k}$ is given by the $k$-th member of (1.2), and the coefficient of $z^{0}$ is unity [13, §3.3]). In the present paper, using the theory of Selberg correlation integrals [16], [2], [20, §4.6], [28], we will show how to systematically obtain a first order $(l+1)\times(l+1)$ matrix linear differential equation, which has as the first component of its vector solution equal to (1.3). As a preliminary step, the matrix average is expressed in terms of a certain Jack polynomial based hypergeometric function of $l$ variables, specialised to the case that all the variables are equal; see Section 2. For a closely related class of hypergeometric functions, earlier work [25] has identified a matrix differential equation characterisation. It is shown in Section 3 how the Jack polynomial based hypergeometric functions of present interest can be deduced as confluent limits of those appearing in [25], allowing for these too to be similarly characterised. This implies a scalar linear differential equation of degree $l+1$ satisfied by (1.3) (Proposition 4.1 below), which with the use of computer algebra can be made explicit. As an illustration, we do this explicitly for $l=8$, which is the first case beyond the listing in [9]; see Proposition 4.2 below.
1.2 A unitary matrix integral and the moments of the second derivative of the Riemann zeta function on the critical line
Motivating taking up the problem of linear differential equation equations satisfied by (1.1) at the present time, and furthermore motivating a more extensive study, is the primary role played by the modification of (1.1) and the Toeplitz determinant in (1.4)
$\Big{\langle}(\det U)^{l}e^{z{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}=\det\Big{[}I_{j-k+l}(2z)\Big{]}_{j,k=0,\dots,l-1}\\ =(-1)^{l(l-1)/2}\det\Big{[}I_{j+k+1}(2z)\Big{]}_{j,k=0,\dots,l-1}$ (1.7)
in the recent studies [35], [36]. Here the second equality views the determinant in the Hankel class (i.e. indices a function of $j+k$) which is the structure emphasised in these studies. The theme of [35], [36] is the question of statistical properties, particularly moments, of higher order derivatives of the Riemann zeta function on the critical line. Fundamentally, in keeping with the celebrated work [34], it is hypothesised that these zeta function moments coincide with the scaled limiting moments of the corresponding derivative of the characteristic polynomial of a Haar distributed random unitary matrix; see [33, 15, 17, 47, 8, 7, 3, 6, 23] for earlier work in this direction, and [4, 44, 1, 5] for subsequent developments. In particular, in relation to the second derivative, this leads to the study of [35]
$F_{2}(l)=\lim_{N\rightarrow\infty}\frac{1}{N^{l^{2}+4l}}\Big{\langle}|Z_{N}^{\prime\prime}(1)|^{2l}\Big{\rangle}_{U(N)},$ (1.8)
where
$Z_{N}(s):=e^{-\pi iN/2}e^{i\sum_{n=1}^{N}\theta_{n}/2}s^{-N/2}\prod_{j=1}^{N}(1-se^{-i\theta_{j}}).$ (1.9)
In [35], it was shown
$F_{2}(l)=(-1)^{\frac{l(l-1)}{2}}\sum_{m=0}^{2l}\binom{2l}{m}\left(\frac{d}{dx}\right)^{4l-2m}\left(e^{-\frac{x}{2}}x^{-m-\frac{l^{2}}{2}}f_{m}(x)\right)\Bigg{|}_{x=0},$ (1.10)
where each $f_{m}(x)$ is a combinatorial linear combination of shifted versions of the Hankel determinant appearing on the left-hand side of (1.7), which takes the form
$\det[I_{j+k+1+2\mu_{j+1}}(2\sqrt{x})]_{j,k=0,\ldots,l-1},$ (1.11)
and where $\mu_{1}+\cdots+\mu_{l}=m$. Moreover, in [36], it was shown to be given in the form
$f_{m}(x)=\frac{1}{x^{m}}\sum_{j=0}^{m}x^{j}P_{j}(x)\frac{d^{j}\tau_{l}(x)}{dx^{j}}.$ (1.12)
Here, each $P_{j}(x)$ is a polynomial defined recursively, and $\tau_{l}(x)$ denotes the Hankel determinant appearing on the left-hand side of (1.7), with $z=2\sqrt{x}$.
In [36], certain reasoning is given that leads to the conclusion that like (1.1) $\tau_{l}(x)$ satisfies a linear differential equation of degree $l+1$ (see Appendix A for an overview). As mentioned in the opening paragraph, our approach to systematically deriving a linear differential equation that (1.1) satisfies allows for an extra factor $(\det U)^{q}$ to be included for general $q$, and thus includes (1.7). Thus in this regards our work complements [36]. We consider this further in §4.3, and discuss too the advantage of the linear differential equation characterisation over an alternative nonlinear differential equation characterisation put forward in [30].
2 Jack polynomial based hypergeometric function form
It has been revised in the Introduction that the matrix integral (1.1) has the interpretations as a partition function for a two-dimensional gauge theory, and as the generating function for counting longest increasing subsequences. There is also an occurrence of (1.1) as a limiting gap probability at the hard edge of a positive definite matrix ensemble with unitary symmetry. To specify this, define complex Wishart matrices $W=X^{\dagger}X$ with $X$ an $n\times N$ $(n\geq N)$ standard complex Gaussian matrix (also referred to as a rectangular GinUE matrix [14]). Let $E_{N}((0,s);a)$ with l=nNl=n-N denote the probability of no eigenvalues in $(0,s)$. In [19] it was established that
$E_{\beta=2}^{\rm hard}((0,s);l):=\lim_{N\to\infty}E_{N}((0,s/4N);l)=e^{-s/4}\Big{\langle}e^{(s^{1/2}/2){\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}.$ (2.1)
The most significant contribution in [19] for present purposes was the identification of the RHS of (2.1) (and generalisations) as a Jack polynomial based generalised hypergeometric function.
To specify this class of multivariable special function, we first recall that, with $\mathbf{x}=(x_{1},\dots,x_{n})$ the Jack polynomials $P_{\kappa}^{(\alpha)}(\mathbf{x})$ are symmetric polynomials of $n$ variables [38, Ch. VI.6], [20, Ch. 12], [37, Ch. 7]. They are dependent on a parameter $\alpha>0$ and labelled by a partition $\kappa:=(\kappa_{1},\dots,\kappa_{n})$; in the latter each part $\kappa_{i}\in\mathbb{Z}_{\geq 0}$ is ordered as $\kappa_{1}\geq\kappa_{2}\geq\cdots\geq\kappa_{n}$. These polynomials are eigenfunctions of the differential operator
$\sum_{j=1}^{n}x_{j}^{2}{\partial^{2}\over\partial x_{j}^{2}}+{2\over\alpha}\sum_{1\leq j<k\leq n}{1\over x_{j}-x_{k}}\Big{(}x_{j}^{2}{\partial\over\partial x_{j}}-x_{k}^{2}{\partial\over\partial x_{k}}\Big{)}.$ (2.2)
Introducing the dominance partial ordering on partitions of the same length $\mu<\kappa$, the Jack polynomial $P_{\kappa}^{(\alpha)}(\mathbf{x})$ is uniquely determined as the eigenfunction of (2.2) having the triangular structure with respect to the monomial basis $\{m_{\mu}(\mathbf{x})\}$,
$P_{\kappa}^{(\alpha)}(\mathbf{x})=m_{\kappa}(\mathbf{x})+\sum_{\mu<\kappa}c_{\kappa,\mu}^{(\alpha)}m_{\mu}(\mathbf{x}).$ (2.3)
In the definition of Jack polynomial based hypergeometric functions, the Jack polynomials play the role of the monomials, although these must be scaled
$P_{\kappa}^{(\alpha)}(\mathbf{x})\mapsto\alpha^{|\kappa|}P_{\kappa}^{(\alpha)}(\mathbf{x})/h_{\kappa}^{\prime},$
where the precise definition of $h_{\kappa}^{\prime}$ (which depends too on $\alpha$) can be found in e.g. [20, Eq. (12.46)]. Playing the role of the Pochhammer symbol is the quantity
$\quad[u]_{\kappa}^{(\alpha)}:=\prod_{j=1}^{n}{\Gamma(u-(j-1)/\alpha+\kappa_{j})\over\Gamma(u-(j-1)/\alpha)};$ (2.4)
see e.g. [20, Eq. (12.46)]. With this notation we then have as the series definition of the generalised hypergeometric function in question (see e.g. [20, §13.1])
${\vphantom{F}}_{p}^{\mathstrut}F_{q}^{(\alpha)}(a_{1},\dots,a_{p};b_{1},\dots,b_{q};\mathbf{x}):=\sum_{\kappa}{\alpha^{|\kappa|}\over h_{\kappa}^{\prime}}\frac{[a_{1}]^{(\alpha)}_{\kappa}\dots[a_{p}]^{(\alpha)}_{\kappa}}{[b_{1}]^{(\alpha)}_{\kappa}\dots[b_{q}]^{(\alpha)}_{\kappa}}P_{\kappa}^{(\alpha)}(\mathbf{x}).$ (2.5)
In the case $n=1$ this is independent of $\alpha$ and reduces to its classical scalar counterpart $_{p}F_{q}$.
We are now in a position to specify the result of [19] which allows the RHS of (2.1) to be identified as a Jack polynomial based generalised hypergeometric function. Thus, for general $\alpha>0$,
${\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(c+(n-1)/\alpha;(t)^{n})=B_{n}(c,\alpha)\Big{(}{1\over t}\Big{)}^{(c-1)n/2}$
$\qquad\times\Big{(}{1\over 2\pi}\Big{)}^{n}\int_{[-\pi,\pi]^{n}}\prod_{j=1}^{n}e^{2t^{1/2}\cos\theta_{j}}e^{i(c-1)\theta_{j}}\prod_{1\leq j<k\leq n}|e^{i\theta_{k}}-e^{i\theta_{j}}|^{2/\alpha}d\theta_{1}\cdots d\theta_{n},$ (2.6)
(see too [20, Eq. (13.27)] and [24, Eq. (4.32)]), where $c\in\mathbb{Z}_{\geq 1}$ (this restriction can be removed by suitably deforming the implied circular contours in (2) to Hankel loops [21, Proof of Prop. 2]) and
$B_{n}(c,\alpha)=\prod_{j=1}^{n}{\Gamma(1+1/\alpha)\Gamma(c+(j-1)/\alpha)\over\Gamma(1+j/\alpha)}.$ (2.7)
The notation $(t)^{n}$ on the LHS denotes that $\mathbf{x}=(t,t,\dots,t)$, and thus all $n$ components are equal to $t$. Making use of (1.5), it follows from (2) that
$\Big{\langle}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(l;(s)^{l})\Big{|}_{\alpha=1};$ (2.8)
for a derivation of this result, and its generalisation for $\alpha=1/2$ and $\alpha=2$, based on the theory of zonal polynomials, see [27]). Extending this, for any $q\in\mathbb{Z}_{\geq 0}$ we read off that
$\Big{\langle}(\det U)^{q}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}={s^{ql/2}\over l!B_{l}(q+1,\alpha)}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(q+l;(s)^{l})\Big{|}_{\alpha=1};$ (2.9)
recall (1.7) in relation to the case $q=l$.
Specifying the circular $\beta$ ensemble CEβ,n is the eigenvalue PDF
${1\over(2\pi)^{n}C_{\beta,n}}\prod_{1\leq j<k\leq n}|e^{i\theta_{k}}-e^{i\theta_{j}}|^{\beta},\quad C_{\beta,n}={\Gamma(\beta n/2+1)\over\Gamma(\beta/2+1)^{n}};$ (2.10)
the case $\beta=2$ corresponds to $U(n)$. This gives the $\beta$ generalisation of (2.9)
$\Big{\langle}(\det U)^{q}e^{s^{1/2}{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{{\rm CE}_{\beta,l}}={s^{ql/2}\over C_{\beta,l}B_{l}(q+1,2/\beta)}{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}(q+1+(l-1)\beta/2;(s)^{l}).$ (2.11)
3 Matrix differential equation characterisation
As in the classical case, there is the limit relation
$\lim_{M\to\infty}{\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(\alpha)}(aM;b;\mathbf{x}/M)={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(\alpha)}(b;\mathbf{x}).$ (3.1)
The significance of this is that for $\mathbf{x}=(t)^{n}$, since the 2022 work [25] we have available a matrix differential equation characterisation of ${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(\alpha)}(a;b;\mathbf{x})$. This comes about in the setting of the integral representation [20, Cor. 2.1]
${\vphantom{F}}_{1}^{\mathstrut}F_{1}^{(2/\beta)}\Big{(}(\beta/2)(n-1)+a+1;\beta(n-1)+a+b+2;(-s)^{n}\Big{)}.\\ ={1\over S_{n}(a,b,\beta)}\int_{0}^{1}dx_{1}\cdots\int_{0}^{1}dx_{n}\,\prod_{l=1}^{n}x_{l}^{a}(1-x_{l})^{b}e^{-sx_{l}}\prod_{1\leq j<k\leq n}|x_{k}-x_{j}|^{\beta},$ (3.2)
valid for Re(a)>1(a)>-1, Re(b)>1(b)>-1. Here $S_{n}(a,b,\beta)$ is the Selberg integral (see e.g. [20, Ch. 4]), which normalises the RHS so that it is equal to unity for $s=0$. The matrix differential equation, which we will first present in the form of a differential-difference recurrence, relates to the family of multi-dimensional integrals
$\hat{H}_{p}(x)={1\over C_{p}^{n}}\int_{0}^{1}dt_{1}\cdots\int_{0}^{1}dt_{n}\,\prod_{l=1}^{n}t_{l}^{a}(1-t_{l})^{b-1}e^{-x\sum_{l=1}^{n}t_{l}}\\ \times\prod_{1\leq j<k\leq n}|t_{k}-t_{j}|^{\beta}e_{p}(1-t_{1},\dots,1-t_{n}),$ (3.3)
for $p=0,1,\dots,n$. The factor $e_{p}(1-t_{1},\dots,1-t_{n})$ in the integrand denotes the $p$-th elementary symmetric polynomials in the variables $\{(1-t_{i})\}_{i=1}^{n}$, while $C_{p}^{n}$ is the binomial coefficient $n$ choose $p$. Setting $p=n$ gives the RHS of (3.2) and thus $_{1}F^{(2/\beta)}_{1}$, up to the Selberg integral normalisation, and setting $p=0$ gives the same with $b$ replaced by b1b-1.
Proposition 3.1.
[25, Cor. 2.5] The multidimesional integrals $\{\hat{H}_{p}(x)\}_{p=0}^{n}$ satisfy the differential-difference system
$(n-p)x\hat{H}_{p+1}(x)=\Big{(}(n-p)x+\tilde{B}_{p}\Big{)}\hat{H}_{p}(x)+x{d\over dx}\hat{H}_{p}(x)-\tilde{D}_{p}\hat{H}_{p-1}(x),\quad(p=0,\dots,n),$ (3.4)
where
$\tilde{B}_{p}=p\Big{(}a+b+1+(\beta/2)(2n-p-1)\Big{)},\qquad\tilde{D}_{p}=p\Big{(}(\beta/2)(n-p)+b\Big{)}.$
Corollary 3.1.
Introduce the column vector $\hat{\mathbf{H}}(x)=[\hat{H}_{p}(x)]_{p=0}^{n}$, and the bi-diagonal matrices
$\mathbf{X}$ $=-{\rm diag}\,[n,n-1,\dots,0]+{\rm diag}^{+}\,[n,n-1,\dots,1]$
$\mathbf{Y}$ $=-{\rm diag}\,[\tilde{B}_{0},\tilde{B}_{1},\dots,\tilde{B}_{n}]+{\rm diag}^{-}\,[\tilde{D}_{1},\tilde{D}_{2},\dots,\tilde{D}_{n}].$ (3.5)
Here $\rm diag$$^{+}$ ($\rm diag$$^{-}$) denotes the leading diagonal above (below) the diagonal itself. We have that the differential-difference equation (3.4) is equivalent to the matrix differential equation
$x{d\over dx}\hat{\mathbf{H}}(x)=(x\mathbf{X}+\mathbf{Y})\hat{\mathbf{H}}(x).$ (3.6)
As a consequence [25, Remark 4.19], introducing the vector of power series solutions
$\hat{\mathbf{H}}(x)=\Big{[}\sum_{k=0}^{\infty}c_{p,k}x^{k}\Big{]}_{p=0}^{n},$
we have
$c_{0,0}=S_{n}(a,b-1,\beta),\quad c_{n,0}=S_{n}(a,b,\beta),\quad c_{n-p,0}=c_{n,0}\prod_{s=1}^{p}{\tilde{B}_{n+1-s}\over\tilde{D}_{n+1-s}},\>\>(p=1,\dots,n)$
and, with $\mathbf{c}_{k}:=[c_{p,k}]_{p=0}^{n}$, the vector recurrence relation
$(k{\mathbf{I}}_{n+1}-\mathbf{Y})\mathbf{c}_{k}=\mathbf{X}\mathbf{c}_{k-1},\>\>\mathbf{c}_{-1}:=\mathbf{0}\>\>(k=1,2,\dots),$ (3.7)
where $\mathbf{I}_{n+1}$ denotes the $(n+1)\times(n+1)$ identity.
We now want to make use of (3.1) in Proposition 3.1, in view of both the first and last components of $\hat{\mathbf{H}}(x)$ relating to $_{1}F^{(2/\beta)}_{1}$.
Proposition 3.2.
Consider the differential-difference system
$(n-p)x\hat{I}_{p+1}(x)=\tilde{B}_{p}\hat{I}_{p}(x)+x{d\over dx}\hat{I}_{p}(x)-p\hat{I}_{p-1}(x),\quad(p=0,\dots,n),$ (3.8)
where $\tilde{B}_{p}$ is as in (3.4). Introduce the column vector $\hat{\mathbf{I}}(x)=[\hat{I}_{p}(x)]_{p=0}^{n}$, and the matrices
$\tilde{\mathbf{X}}={\rm diag}^{+}\,[n,n-1,\dots,1],\quad\tilde{\mathbf{Y}}=-{\rm diag}\,[\tilde{B}_{0},\tilde{B}_{1},\dots,\tilde{B}_{n}]+{\rm diag}^{-}\,[1,2,\dots,n].$ (3.9)
We have that the differential-difference equation (3.8) is equivalent to the matrix differential equation
$x{d\over dx}\hat{\mathbf{I}}(x)=(x\tilde{\mathbf{X}}+\tilde{\mathbf{Y}})\hat{\mathbf{I}}(x).$ (3.10)
Furthermore, introducing the vector of power series solutions
$\hat{\mathbf{I}}(x)=\Big{[}\sum_{k=0}^{\infty}d_{p,k}x^{k}\Big{]}_{p=0}^{n},$ (3.11)
normalised by choosing $d_{0,0}=1$, we have
$\hat{I}_{0}(x)$ $={\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big{(}\beta(n-1)+a+b+1;(x)^{n}\Big{)},$ (3.12)
$\hat{I}_{n}(x)$ $=d_{n,0}\,{\vphantom{F}}_{0}^{\mathstrut}F_{1}^{(2/\beta)}\Big{(}\beta(n-1)+a+b+2;(x)^{n}\Big{)},$ (3.13)
and, provided all the $\tilde{B}_{s}$ are nonzero,
$d_{p,0}={p!\over\prod_{s=1}^{p}\tilde{B}_{s}},\>\>(p=1,\dots,n).$ (3.14)
The coefficient vector $\mathbf{d}_{k}:=[d_{p,k}]_{p=0}^{n}$ satisfies the recurrence relation
$(k{\mathbf{I}}_{n+1}-\tilde{\mathbf{Y}})\mathbf{d}_{k}=\tilde{\mathbf{X}}\mathbf{d}_{k-1},\>\>(k=1,2,\dots).$ (3.15)
Proof.
To deduce (3.8) and (3.10), we substitute in (3.4) and (3.6), respectively, $x\mapsto-x/a$, $\hat{H}_{p}(-x/a)=b^{p}\hat{I}_{p}(x)$ and take the limit $a\to\infty$ with $a+b$ fixed. The evaluations (3.13) follow from the evaluations of $\hat{H}_{n}(x)$, $\hat{H}_{0}(x)$ in terms of $\vphantom{F}{}_{1}^{\mathstrut}F_{1}^{(2/\beta)}$ noted in the sentence above Proposition 3.1. The evaluation (3.14) can be calculated as the eigenvector of $\tilde{\mathbf{Y}}$ corresponding to the eigenvalue 0, normalised to have its first component equal to unity. The vector recurrence relation (3.15) follows by substituting (3.11) in (3.10) and equating like powers of $x$. ∎
[…]
$\tilde{B}_{p}=p\Big{(}q+1+(\beta/2)(l-p)\Big{)},$ (3.16)
as follows from (3.13). This for $q\in\mathbb{Z}_{\geq 0}$ and $p=0,\dots,l$ is non-negative, allowing then for the applicability of (3.15) written in the form
$\mathbf{d}_{k}=(k{\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}\mathbf{d}_{k-1},\>\>(k=1,2,\dots)$ (3.17)
and with the initial condition $\mathbf{d}_{0}=[d_{p,0}]_{p=0,\dots,n}$, where $d_{0,0}=1$ and $d_{p,0}$ ($p\geq 1$) is as given by (3.14). Further specialising to $q=0$, $\beta=2$, we can use (3.17) to compute $\{T_{l}(N)\}_{N=1,,\dots,N^{*}}$ (here $N^{*}$ is some given largest value of $N$) according to the matrix product formula
$T_{l}(N)=(N!)^{2}(\mathbf{R}_{N,l}\,\mathbf{d}_{0})_{1},$ (3.18)
where the notation $(\cdot)_{1}$ denotes the first entry in the corresponding column vector and
$\mathbf{R}_{N,l}:=(N{\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}((N-1){\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}\cdots({\mathbf{I}}_{l+1}-\tilde{\mathbf{Y}})^{-1}\tilde{\mathbf{X}}$.
As a concrete example, implementation of (3.18) with $l=4$, and thus involving products of $5\times 5$ matrices, allows for generation of the sequence (1.2).
Remark 3.1.
1. Suppose we replace $q$ by q1q-1 in (3.16). Then setting $q=0$ we observe that the column vector $\delta_{p,l}]_{p=0}^{l}$ is an eigenvector of $\tilde{\mathbf{Y}}$ with $n=l$, allowing for the choice of initial vector corresponding to $d_{p,0}=\delta_{p,l}$ ($p=0,\dots,l$). Using this initial vector, an alternative formula to (3.18) is
$T_{l}(N)=(N!)^{2}(\mathbf{R}_{N,l}\,\mathbf{d}_{0})_{l+1},$ (3.19)
where the notation $(\cdot)_{l+1}$ denotes the $(l+1)$-th entry (final entry) in the corresponding column vector. In the following section, when referring to $\hat{I}_{n}(x)$ in the case $q=0$ as it relates to (2.8), it will be assumed that the initial vector has been similarly modified (and thus in (3.13) $d_{n,0}$ is replaced by unity — without this modification (3.13) is invalid since $d_{n,0}$ involves a division by zero).
2. For efficient computation using the matrix differential equation, it is advantageous to forgo (3.18) in favour of a recursive computation based on (3.15) — we take up this point further at the end of §4.2.
4 Scalar differential equation
4.1 The case of general parameters
Next we would like to relate (3.10) to a scalar linear differential equation of degree $n+1$ for $\hat{I}_{n}(x)$ (and thus by the replacement $b\mapsto b-1$ for $\hat{I}_{0}(x)$; recall (3.12)).
Proposition 4.1.
Let $\hat{I}_{n}(x)$ be specified via the differential-difference system (3.8). For some polynomials $\{q_{m}(x)\}_{m=0,1,\dots,n-1}$ we have
$\Big{(}x^{n}{d^{n+1}\over dx^{n+1}}+q_{n}(x)x^{n-1}{d^{n}\over dx^{n}}+q_{n-1}(x)x^{n-2}{d^{n-1}\over dx^{n-1}}+\cdots+q_{2}(x)x{d^{2}\over dx^{2}}+q_{1}(x){d\over dx}+q_{0}(x)\Big{)}\hat{I}_{n}(x)=0.$ (4.1)
The degree in $x$ of the coefficient of each $d^{j}\over dx^{j}$} is [(n1+j)/2][(n-1+j)/2], thus restricting the degrees of the $q_{m}(x)$. Furthermore, when $\{\tilde{B}_{j}\}$ are all integers, the coefficients of these polynomials are also integers.
Proof.
Let $D_{x}:=x{d\over dx}$. From (3.8) in the case $p=n$ we read off that
$n\hat{I}_{n-1}(x)=(\tilde{B}_{n}+D_{x})\hat{I}_{n}(x).$ (4.2)
Using this in the case p=n1p=n-1 of (3.8) then gives
$n(n-1)\hat{I}_{n-2}(x)=\Big{(}(\tilde{B}_{n-1}+D_{x})(\tilde{B}_{n}+D_{x})-nx\Big{)}\hat{I}_{n}(x).$ (4.3)
Using both (4.2) and (4.3) in the case p=n2p=n-2 of (3.8) then gives
$n(n-1)(n-2)\hat{I}_{n-3}(x)=\bigg{(}(\tilde{B}_{n-2}+D_{x})\Big{(}(\tilde{B}_{n-1}+D_{x})(\tilde{B}_{n}+D_{x})-nx\Big{)}-2x(n-1)(\tilde{B}_{n}+D_{x})\bigg{)}\hat{I}_{n}(x).$ (4.4)
We see that the differential operators on the RHS of (4.2), (4.3), (4.4) have the structure
$x{d\over dx}+c_{0},\>\>x^{2}{d^{2}\over dx^{2}}+d_{0}x{d\over dx}+(c_{1}+c_{2}x),\>\>x^{3}{d^{3}\over dx^{3}}+e_{0}x^{2}{d^{2}\over dx^{2}}+(d_{1}+d_{2}x)x{d\over dx}+(c_{3}+c_{4}x).$ (4.5)
Repeating this procedure down to the case $p=1$ of (3.8) allows each $\hat{I}_{j}(x)$ ($j=0,\dots,n)$ to be written as a differential operator of degree njn-j acting on $\hat{I}_{n}(x)$. In particular, $n!\hat{I}_{1}(x)$ has the structure of a differential operator of the form (4.1) with $n\mapsto n-1$ and an extra factor of $x$ multiplying each of the terms involving a derivative. Similarly, $n!\hat{I}_{0}(x)$ has the structure of a differential operator of the form in (4.1) as written, except for an extra factor of $x$ multiplying each of the terms involving a derivative.
The case $p=0$ of (3.8) gives
${d\over dx}\hat{I}_{0}(x)-n\hat{I}_{1}(x)=0.$
From the established formulas for $\hat{I}_{0}(x),\hat{I}_{1}(x)$, this implies (4.1). That the coefficients of the polynomials $\{q_{j}(x)\}$ are integers for $\{\tilde{B}_{j}\}$ integers is immediate from the structure of the RHS of (4.2)–(4.4) as extends to $\hat{I}_{0}(x),\hat{I}_{1}(x)$. ∎
Remark 4.1.
The restriction on the degrees of the polynomials $\{q_{j}(x)\}$ relates to a conjecture from [10]. See too the discussion in [9].
4.2 Characterisation of the matrix integral in (1.3)
In the case $\beta=2$, $n=l$ and $\tilde{B}_{p}=p(l-p)$, $\hat{I}_{n}(x)$ is according to (2.9) and (3.13) (the latter modified in keeping with Remark 3.1.1) equal to the matrix average in (1.3) with $z=x^{1/2}$ and thus is a generating function for the sequence $\{T_{l}(N)\}_{N=1,2,\dots}$. Using different methods, the precise form of the corresponding differential operator in (4.1) has been computed for $l=2,3,\dots,7$ in [9]. From the substitution method given in the proof of Proposition 3.1 we can reclaim the listing of [9], and extend it to higher values of $l$ (see the work [41] for other applications in random matrix theory of this method in reducing first order linear matrix differential equations to higher order scalar differential equations). Here we make note of the case $l=8$.
Proposition 4.2.
The matrix integral in (1.3) with $z=x^{1/2}$ and $l=8$ satisfies the differential equation in (4.1) with $n=8$ (and thus is of degree 9), with
$q_{8}(x)=120,\>\>\>q_{7}(x)=3(1477-40x),\>\>\>q_{6}(x)=2(48809-4467x),$
$q_{5}(x)=6(203139-42433x+728x^{2}),\>\>\>q_{4}(x)=4(2177225-881163x+51840x^{2}),$
$q_{3}(x)=4(8411053-6187600x+856062x^{2}-13120x^{3}),$
$q_{2}(x)=(7322954-10406303x+2993572x^{2}-161622x^{3}),$
$q_{1}(x)=16(1456832-6870928x+4305495x^{2}-584784x^{3}+9216x^{4}),$
$q_{0}(x)=32(-728416+1821603x-643914x^{2}+36864x^{3}).$
In relation to this characterisation for general $l$, substituting the series on the RHS of (1.3) transforms the scalar linear differential equation to a scalar linear difference equation with coefficients that are polynomials in $N$, and of degree no greater than $l+1$; see [9] for more on this and [11, footnote 22] for references containing the explicit forms up to $l=5$. However, in relation to computing the power series solution, there is no advantage in first reducing the matrix differential equation down to a scalar differential equation, rather use of the analogue of (3.18) (which is based on the matrix differential equation) is more efficient.
On this last point, it is of interest to compare our scheme for the computation of $\{T_{l}(N)\}$, which we will base on (3.15), to that used in the recent works [26, 11]. There, instead of considering (1.1) directly, one first aims to provide a power series expansion of what is essentially the logarithmic (with $z={1\over 2}\sqrt{s}$),
$u(s):=-s{d\over ds}\log\bigg{(}e^{-s/4}\Big{\langle}e^{{1\over 2}\sqrt{s}{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}\bigg{)}.$ (4.6)
The significance of (4.6) is the characterisation of $u(s)$ as the solution of the $\sigma$-Painlevé III equation [45, 29]
$(su^{\prime\prime})^{2}+u^{\prime}(4u^{\prime}-1)(u-su^{\prime})=0$ (4.7)
with boundary condition for small $s$
$u(s)=-{s^{1+l}\over 2^{2(1+l)}\Gamma(1+l)\Gamma(2+l)}+{\rm O}(s^{2+l}).$ (4.8)
Writing $u(s)$ as a power series consistent with (4.8) and substituting in (4.7) provides a recurrence for successive coefficients. In this way, after integrating and exponentiating as required by (4.6), we have available terms in the power series expansion of the average in (4.6), and thus from (1.3) (after doing this for $l=1,2,\dots,N^{*}$) knowledge of $\{T_{l}(N)\}_{l=1,\dots,N}$ for values of $N$ up to $N^{*}$. By carrying out this procedure, tabulation up to $N^{*}=700$ was obtained in [26]. A computational improvement of this strategy was provided in [11], based on the observation that differentiating (4.7) to convert it to a particular third order nonlinear equation in the Chazy-I class. The latter is linear in the higher derivative ((4.7) is quadratic), and is quadratic in lower order terms ((4.7) is cubic). Based on this third order nonlinear equation, a tabulation up to $N^{*}=1000$ was attained in [11]. On the other hand, the recurrence (3.15) of Proposition 3.2 provides an alternative to the use of a nonlinear equation for purposes of computing the tabulation. In this formalism $T_{l}(N^{*})$ is read off from the vector $\mathbf{d}_{N^{*}}$. With the vector $\mathbf{d}_{0}$ known from (3.14), successive vectors are computed by solving the linear system in the recurrence (3.15) by back substitution. This only takes of order $l$ steps since the matrix on the right hand side is bidiagonal. With $N^{*}$ iterations the total number of steps in then of order $N^{*}l$, which for small $l$ at least thus provides a computational saving relative to the method of [11]. One remarks that the use of such tabulations was to probe correction terms to the leading scaled functional form of the distribution of the length of the longest increasing subsequence [26, 11, 12].
4.3 Characterisation of the matrix integral in (1.7)
The result of Proposition 3.2 with $\beta=2$, $n=l$, $\tilde{B}_{p}=p(2l-p)$ gives, according to (2.9) and (3.13), the matrix integral in (1.7) as proportional to $\hat{I}_{n}(x)$. Following the method of Proposition 4.2, we can similarly specify the polynomials in the scalar linear differential equation reduction of the matrix differential equation from Proposition 3.2 for small values of $l$ (without this latter restriction, the coefficients in the corresponding polynomials soon become unwieldy due to the number of digits involved and their increasing degree; recall the listing of Proposition 4.2).
Proposition 4.3.
Consider the matrix integral in (1.7) with $z=x^{1/2}$, multiplied by $x^{-l^{2}/2}$ and up to a sign factor $(-1)^{\frac{l(l-1)}{2}}$. For $l=2,\dots,5$ this quantity satisfies the linear differential equations
$\bigg{(}x^{2}{d^{3}\over dx^{3}}+10x{d^{2}\over dx^{2}}+(20-4x){d\over dx}-10\bigg{)}y(x)=0,$
$\bigg{(}x^{3}{d^{4}\over dx^{4}}+28x^{2}{d^{3}\over dx^{3}}+(230-10x)x{d^{2}\over dx^{2}}+(540-122x){d\over dx}+(-270+9x)\bigg{)}y(x)=0$
$\bigg{(}x^{4}{d^{5}\over dx^{5}}+60x^{3}{d^{4}\over dx^{4}}+(1238-20x)x^{2}{d^{3}\over dx^{3}}+(10268-662x)x{d^{2}\over dx^{2}}+(28288-6108x+64x^{2}){d\over dx}$
${}\qquad\qquad+(-14144+640x)\bigg{)}y(x)=0$
$\bigg{(}x^{5}{d^{6}\over dx^{6}}+110x^{4}{d^{5}\over dx^{5}}+(4538-35x)x^{3}{d^{4}\over dx^{4}}+(87008-2422x)x^{2}{d^{3}\over dx^{3}}+(765908-55800x+$
${}\quad+259x^{2})x{d^{2}\over dx^{2}}+(2431000-485408x+8392x^{2}){d\over dx}+(-1215500+59300x-225x^{2})\bigg{)}y(x)=0,$
respectively.
As revised in Appendix A, the work [36] has given an alternative way to derive the degree $l+1$ scalar linear differential equation satisfied by the matrix integral (1.7). In the case $l=2$ this is made explicit in [36, Eq. (97)], and consistency with the third order equation listed above can be checked. Implementing the recursive strategy leading to (A.12) of the Appendix in the cases $l=3,4,5$ using computer algebra establishes consistency of the results obtained by the alternative method for all the differential equations listed in Proposition 4.3.
Motivating the investigation of [36] in this direction was the realisation that the formalism given in [30] relating to a characterisation in terms of a Painlevé III $\sigma$-form transcendent cannot be used to generate the full power series about the origin due to the stated boundary condition therein not uniquely determining the coefficients. Thus, considered in [30] was the quantity
$v(s):=-s{d\over ds}\log\bigg{(}s^{l^{2}/2}e^{-s/4}\Big{\langle}(\det U)^{l}e^{{1\over 2}\sqrt{s}{\rm Tr}(U+U^{\dagger})}\Big{\rangle}_{U(l)}\bigg{)};$ (4.9)
cf. (4.6). Using results from [29] it was noted that $v(s)$ satisfies the particular $\sigma$-Painlevé III equation
$(sv^{\prime\prime})^{2}+v^{\prime}(4v^{\prime}-1)(v-sv^{\prime})-{l^{2}\over 16}=0,$ (4.10)
(cf. (4.7)) subject to the small-$s$ boundary condition
$v(s)=-l^{2}+{s\over 8}+{\rm O}(s^{2}).$ (4.11)
It was shown in [30] that this allowed for the generation of the first $l$ even powers recursively, as was the need of the application therein. However, missed in [30] was the fact made explicit in [36] that the boundary condition (4.11) does not uniquely determine a solution of (4.11), with their in fact being a family of solutions which require the specification of the coefficient of $s^{2l+1}$ to be made unique beyond degree $s^{2l}$, and which was incorrectly set to zero.
The finding in [36] of a scalar differetial equation characterisation of the matrix integral (1.7) overcame this latter point, and was used to obtain in particular the coefficients up to degree $s^{4l}$ as required for the application to the moments of the second derivative of the Riemann zeta function on the critical line as outlined in the text of the Introduction including (1.8). The present work complements such a computational objective, as in keeping with the text below (4.7) use can be made of the matrix differential equation formalism to provide efficient tabulations based on solving a vector recurrence via back substitution.
Acknowledgement
[…]
Appendix A
The purpose of this Appendix is to outline the method given in [36] for the derivation of the degree $l+1$ linear differential equation satisfied by (1.7) (see Corollary A.1 below), which at a practical level allows for an alternative derivation of the results of Proposition 4.3. Moreover, the working of [36] will be extended to allow for consideration of the family of matrix averages in (2.9) with $q=l+\alpha$ (for $q\in\mathbb{Z}_{\geq 0}$), or more generally the family of Hankel determinants in (1.7) with the index of the $I$-Bessel function therein generalised from $j+k+1$ to $j+k+\alpha$ (for $\alpha\in\mathbb{R}_{\geq 0}$). The starting point of the method is the connection between a shifted version of the Hankel determinant appearing in (1.7)
$\tau_{l}(x):=\det\!\big{[}I_{j+k+1}(2\sqrt{x})\big{]}_{j,k=0}^{\,l-1},$ (A.1)
and the derivative of the Hankel determinant itself. Specifically, we begin with the definition of the determinant of Hankel matrices whose column indices are shifted by Young diagrams.
Definition 1.
Let $\alpha\in\mathbb{R}_{\geq 0}$. Let $l\geq 1,h\geq 0$ be integers. Let $Y=(m_{1},\ldots,m_{h})$ be a Young diagram. Define
$\tau_{l,Y,\alpha}(x):=\det\!\big{[}I_{j+k+\alpha+m_{l-k}}(2\sqrt{x})\big{]}_{j,k=0}^{\,l-1}.$
Here we adopt the convention that $m_{h+1}=\cdots=m_{l}=0$ whenever $h<l$. When $Y=\emptyset$, we simply write $\tau_{l,\emptyset,\alpha}$ as $\tau_{l,\alpha}$ for short.
In particular, when $Y=\emptyset$ and $\alpha=1$, the above definition reduces to $\tau_{l}(x)$ as given in (A.1). Using the trace of a square matrix, we define the following operations.
Definition 2.
Let $\alpha\in\mathbb{R}_{\geq 0}$. Let $l\geq 1$, $h,n\geq 0$ be integers. Let $Y=(m_{1},\ldots,m_{h})$ be a Young diagram. Define
$(T_{n}\tau_{l,Y,\alpha})(x):=\mathrm{Tr}\left(\mathrm{adj}\big{(}\big{[}I_{j+k+\alpha+m_{l-k}}(2\sqrt{x}\big{)}\big{]}_{j,k=0}^{\,l-1})\big{[}I_{j+k+\alpha+m_{l-k}+n}(2\sqrt{x})\big{]}_{j,k=0}^{\,l-1}\right)$
and
$(S_{n}\tau_{l,Y,\alpha})(x):$
$=\mathrm{Tr}\left(\mathrm{adj}\big{(}\big{[}I_{j+k+\alpha+m_{l-k}}(2\sqrt{x})\big{]}_{j,k=0}^{\,l-1}\big{)}\big{[}(j+k+\alpha+m_{l-k}+n)I_{j+k+\alpha+m_{l-k}+n}(2\sqrt{x})\big{]}_{j,k=0}^{\,l-1}\right),$
where for a matrix $M$, $\mathrm{adj}(M)$ denotes its adjugate.
In deriving the higher order linear differential equation for $\tau_{l,\alpha}$, we mainly use a special form of Young diagram, namely
$Y_{m,j}:=(m-j+1,1,1,\ldots,1),$ (A.2)
of weight $m$, where $m\geq j\geq 1$. By applying [36, Thm. 3.1] to general Hankel determinants shifted by $Y_{m,j}$ and specializing to $\tau_{l,Y_{m,j},\alpha}$, we obtain the following inverse linear system.
Proposition A.4.
Let $l,m\geq 1,m\geq j\geq 1$. Let $Y_{m,j}$ be given as in (A.2). Then
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\\ \end{pmatrix}=B^{(m)}\begin{pmatrix}\vdots\\ \sum_{h=1}^{j-1}(-1)^{h}T_{h}\tau_{l,Y_{m-h,j-h},\alpha}\\ \vdots\\ T_{m}\tau_{l,\alpha}\\ \end{pmatrix}_{j=2,\ldots,m},$
where $B^{(m)}=(b_{j,k}^{(m)})_{i,j=1,\ldots,m}$ is the $m\times m$-matrix satisfying
$b_{j,k}=\begin{cases}(-1)^{j+k-1}/k(k+1)&{\color[rgb]{0,0,0}j\leq k\leq m-1};\\ -1/j&k=j-1;\\ 0&k<j-1;\\ (-1)^{j-1}/m&k=m.\end{cases}.$ (A.3)
Using the recurrence relations of the modified Bessel function of the first kind,
$\frac{d}{dx}I_{\beta}(2\sqrt{x})=\frac{I_{\beta+1}(2\sqrt{x})}{\sqrt{x}}+\frac{\beta}{2x}I_{\beta}(2\sqrt{x}),\quad I_{\beta+2}(2\sqrt{x})=I_{\beta}(2\sqrt{x})-\frac{\beta+1}{\sqrt{x}}I_{\beta+1}(2\sqrt{x}),$ (A.4)
we can express $\tau_{l,Y_{1,1},\alpha}$, $\tau_{l,Y_{2,1},\alpha}$, and $\tau_{l,Y_{2,2},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. The following lemmas for general $\alpha$ are obtained by a similar argument to that used in [36, Prop. 4.1 and Prop. 4.2] for $\alpha=1$.
Lemma 1.
Let $l\geq 1,m\geq 0$ and $Y$ be a Young diagram of length $m$. Then
$T_{1}\tau_{l,Y,\alpha}(x)=\sqrt{x}\frac{d}{dx}\tau_{l,Y,\alpha}(x)-\frac{l^{2}+m+l(\alpha-1)}{2\sqrt{x}}\tau_{l,Y,\alpha}.$
In particular, if $Y=\emptyset$, then
$\tau_{l,Y_{1,1},\alpha}(x)=\sqrt{x}\frac{d}{dx}\tau_{l,\alpha}(x)-\frac{l^{2}+l(\alpha-1)}{2\sqrt{x}}\tau_{l,\alpha}.$ (A.5)
Lemma 2.
Let $l\geq 1$. Then
$\tau_{l,Y_{2,1},\alpha}=$ $\frac{x}{2}\frac{d^{2}}{dx^{2}}\tau_{l,\alpha}-\frac{l(l+\alpha+1)+\alpha-1}{2}\frac{d}{dx}\tau_{l,\alpha}$
$+\frac{(l^{2}+l(\alpha-1))(l^{2}+l(\alpha+3)+2\alpha)+4lx}{8x}\tau_{l,\alpha},$ (A.6)
$\tau_{l,Y_{2,2},\alpha}=$ $\frac{x}{2}\frac{d^{2}}{dx^{2}}\tau_{l,\alpha}-\frac{l(l+\alpha-3)-(\alpha-1)}{2}\frac{d}{dx}\tau_{l,\alpha}$
$+\frac{(l^{2}+l(\alpha-1))(l^{2}+l(\alpha-5)+4-2\alpha)-4lx}{8x}\tau_{l,\alpha}.$ (A.7)
For a general weight $m$ of Young diagrams, we can recursively obtain an expression for $\tau_{l,Y_{m,j},\alpha}$ in terms of the derivatives of $\tau_{l,\alpha}$. Thus, by identity (A.4), we can rewrite operations involving $T_{h}$ in Proposition A.4 in terms of $S_{h-1}$ and $T_{h-2}$. Combining Lemmas 1 and 2, we then obtain the following formula for $\tau_{l,Y_{m,j},\alpha}$, which depends recursively on $m$.
Proposition A.5.
Let $l\geq 1$ and $m\geq 3$. Let $m\geq j\geq 1$. Let $Y_{m,j}$ be given as in (A.2). Then
$\begin{pmatrix}\tau_{l,Y_{m,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m,m},\alpha}\end{pmatrix}$= $-\,\sqrt{x}B^{(m)}\begin{pmatrix}\frac{d}{dx}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \frac{d}{dx}\tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}+\frac{l^{2}+l(\alpha-1)+m-1}{2\sqrt{x}}B^{(m)}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\\ 0\end{pmatrix}$
$-\,\frac{1}{\sqrt{x}}C_{1}^{(m)}\begin{pmatrix}\tau_{l,Y_{m-1,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-1,m-1},\alpha}\end{pmatrix}+C_{2}^{(m)}\begin{pmatrix}\tau_{l,Y_{m-2,1},\alpha}\\ \vdots\\ \tau_{l,Y_{m-2,m-2},\alpha}\end{pmatrix}$, (A.8)
where $B_{1}^{(m)}$ is given as in (A.3), $C_{1}^{(m)}=(c_{j,k}^{(1)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-1\end{subarray}}$ be an $m\times(m-1)$-matrix satisfying
$c_{j,k}^{(1)}=\begin{cases}(-1)^{j+k}\big{(}\frac{m-1-k}{k+1}+\frac{2l-k+\alpha}{k(k+1)}\big{)}&\text{if }j\leq k\leq m-1;\\ -\frac{1}{k+1}\big{(}k(2l-k+\alpha)+m-1-k\big{)}&\text{if }k=j-1;\\ 0&\text{if }k<j-1,\end{cases},$ (A.9)
and $C_{2}^{(m)}=(c_{j,k}^{(2)})_{\begin{subarray}{c}j=1,\ldots,m\\ k=1,\ldots,m-2\end{subarray}}$ is an $m\times(m-2)$-matrix satisfying
$c_{j,k}^{(2)}=\begin{cases}(-1)^{j+k}\frac{l+2}{(k+1)(k+2)}&j-1\leq k\leq m-2;\\ \frac{j-l-2}{j}&k=j-2;\\ 0&k<j-2.\end{cases}.$ (A.10)
Remark A.2.
The above proposition generalizes [36, Prop. 4.3] from the case $\alpha=1$ to general $\alpha$. It can be rewritten as a first order matrix linear differential equation for the vector on the LHS analogous to (but distinct from) that implied by Proposition 3.2, with $n=m$, $\beta=2$ and $\tilde{B}_{p}$ given by (3.16) with $q=m+\alpha$. On this latter point, note that the entries in the corresponding vectors agree in the final entry but involve different functions otherwise.
Theorem 1.
Let $l\geq 1$. Suppose that $\tau_{l,Y_{l+1,l+1},\alpha}$ is obtained recursively by formula (A.5) with initial conditions (A.5), (A.6), and (A.7). Then
$\tau_{l,Y_{l+1,l+1},\alpha}(x)=x^{-\frac{l+1}{2}}\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{\lfloor\frac{l+1-n}{2}\rfloor}b_{j,n}^{(l+1)}(\alpha)\,x^{j}\right)\frac{d^{n}\tau_{l,\alpha}(x)}{dx^{n}},$ (A.11)
where the coefficients $b_{j,n}^{(l+1)}(\alpha)$ are given recursively. Moreover, we have the following equation
$\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{\lfloor\frac{l+1-n}{2}\rfloor}b_{j,n}^{(l+1)}(\alpha)\,x^{j}\right)\frac{d^{n}\tau_{l,\alpha}(x)}{dx^{n}}=0.$ (A.12)
Proof.
By induction, it is straightforward to verify that (A.11) holds. It now remains to prove (A.12). On one hand, by Proposition A.5,
$\tau_{l,Y_{l+1,l+1},\alpha}(x)=$ $\frac{1}{l+1}\left(\sqrt{x}\frac{d\tau_{l,Y_{l,l},\alpha}(x)}{dx}-\frac{l^{2}+l\alpha}{2\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)\right)+\frac{l(l+\alpha)}{(l+1)\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)$
$-\frac{1}{l+1}\tau_{l,Y_{l-1,l-1},\alpha}(x).$ (A.13)
On the other hand, note that
$\tau_{l,Y_{l-1,l-1},\alpha}(x)=\mathrm{Tr}\left(\mathrm{adj}\big{(}\big{[}I_{j+k+\alpha+1}(2\sqrt{x}\big{)}\big{]}_{j,k=0}^{\,l-1})\big{[}I_{j+k+\alpha}(2\sqrt{x})\big{]}_{j,k=0}^{\,l-1}\right).$
By the observation $\tau_{l,Y_{l,l},\alpha}(x)=\det(I_{j+k+\alpha+1}(2\sqrt{x}))_{j,k=0,\ldots,l-1}$ and the recurrence relation
$I_{j+k+\alpha}(2\sqrt{x})=\sqrt{x}\frac{d}{dx}I_{j+k+\alpha+1}(2\sqrt{x})+\frac{\alpha+j+k+1}{2\sqrt{x}}I_{j+k+\alpha+1}(2\sqrt{x}),$
we have
$\sqrt{x}\frac{d}{dx}\tau_{l,Y_{l,l},\alpha}(x)+\frac{l^{2}+l\alpha}{2\sqrt{x}}\tau_{l,Y_{l,l},\alpha}(x)-\tau_{l,Y_{l-1,l-1},\alpha}(x)=0.$
Then by (A.13), we have the identity
$\tau_{l,Y_{l+1,l+1},\alpha}\equiv 0.$
Hence, using above and (A.11), we obtain (A.12). ∎
Remark A.3.
Not only $\tau_{l,Y_{l+1,l+1},\alpha}$, but also $\tau_{l,Y_{m,j},\alpha}$ for any $m,j$ with $l+1\leq j\leq m$, which can be obtained recursively from formula (A.5) as explained in [36, Remark 3.5], should be 0. Hence, this will also lead to a linear differential equation of a form similar to (A.12), but of degree $m$.
As a consequence of Theorem 1 with $\alpha=1$, we obtain the following result, which provides a method to derive the degree $l+1$ linear differential equation satisfied by (1.7).
Corollary A.1.
Let $l\geq 1$. Let $b_{j,n}^{(l+1)}(1)$ be given as in Theorem 1 with $\alpha=1$. Then
$\sum_{n=0}^{l+1}x^{n}\left(\sum_{j=0}^{\lfloor\frac{l+1-n}{2}\rfloor}b_{j,n}^{(l+1)}(1)\,x^{j}\right)\frac{d^{n}\tau_{l}(x)}{dx^{n}}=0.$
Finally, applying Corollary A.1 and verifying with the code [^1] https://drive.google.com/file/d/1gd4kQO9ABKhALrzeYos-DbeLcelQmik6/view?usp=share_link, we re-obtain Proposition 4.3.